OpenStudy (anonymous):

how do i graph y=1/5x?

OpenStudy (anonymous):

just apply any value of x and you'll get the y counterpart

OpenStudy (anonymous):

start at (0,0) go up 1 then over 5 (5,1)

OpenStudy (anonymous):

can you give an example?

OpenStudy (anonymous):

let x be 1 then y=1/5

OpenStudy (anonymous):

your slope is 1/5=rise/run